Grade 10Mechanics

SHM:Period of particle performing SHM is 8s At t=0 particle is at mean position Ratio of distances covered in 1st-second to the 2nd-second is?(I found it to be (1:√2)But the answer is √2+1|HOW?)

Norman Manchu
8 Years agoGrade 10

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
 let the equation of motion is X =Asin(2px/8) in first second X1 =Asin(p/4) in 2 sec X2 =Asin(p/2) so 2nd sec distance covered = X2 -X1 x1 /(X2 -X1) = Asin(p/4)/[Asin(p/2) -Asin(p/4)] =1/(sqrt(2) -1)
 = [sqrt(2) + 1 ] : 1